cs161: Data Structures and Algorithms Handout #2 Michael Goldwasser Stanford University Tuesday, June 25, 1996

Sign-up Questionnaire

In order to shape the course for the audience, we hope to get a good picture of the experience of the students in the class. Please fill out the questionnaire below and return it to us. Feel free to give any explanations that will help.

Name: E-mail: Do you have access to the World Wide Web? (we will put postscript versions of handouts there for viewing/printing) Major/Program: Year: Why are you taking this course?

Related courses you have taken: What programming experience do you have? In what language(s)? What data structures have you seen previously? (e.g. linked lists, binary trees, priority queues)

What algorithms have you seen previously? (e.g. quicksort, binary search, minimum spanning tree algorithms)